What is the ratio of the escape speed of Mars to that of the Earth, ( Mearth = 9.31*Mmars; Rearth=1.88*RMars)

A) 0.23                       (B) 0.45                               (C) 0.66                       (D) 0.20
Answer #1

vearth = sqrt(2GM/r)

M(mars) = (1/9.31) * M(earth) = 0.107 M(earth)

r(mars) = (1/1.88) * r(earth) = 0.531 r(earth)

vmars = sqrt(2GM(0.105)/[r*0.531])

vmars = sqrt(0.105/0.531) * vearth

=> vmars = sqrt(0.105/0.531) * vearth

=> vmars = 0.45 * vearth

Hence, OPTION B
Answer #2

(B) 0.45   
Answer #3

correct answer is

(B) 0.45
Answer #4

(B) 0.45